Given f(0) = – 18, f(1) = 0, f(3) = 0, f(5) = – 248, f(6) = 0, f(9) = 13104; find f(x).


1
Expert's answer
2021-01-12T00:40:38-0500
"\begin{matrix}
 x_0=0 & y_0=-18\ \ \\
 x_1=1 & y_1=0\ \ \ \ \ \ \ \\
x_2=3 & y_2=0\ \ \ \ \ \ \ \\
x_3=5 & y_3=-248\\
x_4=6 & y_4=0\ \ \ \ \ \ \ \\
x_5=9 & y_5=13104\\
\end{matrix}"


The interpolating polynomial is:

"L(x)=\dfrac{(x-x_1)(x-x_2)(x-x_3)(x-x_4)(x-x_5)}{(x_0-x_1)(x_0-x_2)(x_0-x_3)(x_0-x_4)(x_0-x_5)}\times y_0"

"+\dfrac{(x-x_0)(x-x_2)(x-x_3)(x-x_4)(x-x_5)}{(x_1-x_0)(x_1-x_2)(x_1-x_3)(x_1-x_4)(x_1-x_5)}\times y_1"

"+\dfrac{(x-x_0)(x-x_1)(x-x_3)(x-x_4)(x-x_5)}{(x_2-x_0)(x_2-x_1)(x_2-x_3)(x_2-x_4)(x_2-x_5)}\times y_2"


"+\dfrac{(x-x_0)(x-x_1)(x-x_2)(x-x_4)(x-x_5)}{(x_3-x_0)(x_3-x_1)(x_3-x_2)(x_3-x_4)(x_3-x_5)}\times y_3"


"+\dfrac{(x-x_0)(x-x_1)(x-x_2)(x-x_3)(x-x_5)}{(x_4-x_0)(x_4-x_1)(x_4-x_2)(x_4-x_3)(x_4-x_5)}\times y_4"

"+\dfrac{(x-x_0)(x-x_1)(x-x_2)(x-x_3)(x-x_4)}{(x_5-x_0)(x_5-x_1)(x_5-x_2)(x_5-x_3)(x_5-x_4)}\times y_5"



"L(x)=\dfrac{(x-1)(x-3)(x-5)(x-6)(x-9)}{(0-1)(0-3)(0-5)(0-6)(0-9)}\times (-18)"

"+0"

"+0"

"+\dfrac{(x-0)(x-1)(x-3)(x-6)(x-9)}{(5-0)(5-1)(5-3)(5-6)(5-9)}\times (-248)"

"+0"

"+\dfrac{(x-0)(x-1)(x-3)(x-5)(x-6)}{(9-0)(9-1)(9-3)(9-5)(9-6)}\times 13104"

"L(x)=(x^2+x+1)(x-1)(x-3)(x-6)"